OpenStudy (anonymous):

the sum of two numbers is 5 times their difference. if one exceeds the other by one, what are the numbers??

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Okay, word problems... first step is to represent your values... but you have to read your problem thoroughly... Now, you have two numbers... say we represent one of them with "x" How do you represent the other number?

OpenStudy (anonymous):

y

terenzreignz (terenzreignz):

I guess you could do that, but that's not the quickest way... Remember, your problem says "one exceeds the other by one" which means your other number is "one more" than your previous number.

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

?

terenzreignz (terenzreignz):

Whoa there... where did you get that?!

OpenStudy (anonymous):

idk like at all lol i just woke up lol so im not workin that great so represent it with maybe 1 ?

OpenStudy (anonymous):

or x+1

terenzreignz (terenzreignz):

x+1, yes :D

OpenStudy (anonymous):

okay x+1 so now what

terenzreignz (terenzreignz):

So we have x, and x + 1, these are our two numbers... So it says... the sum of these two numbers is equal to five times their difference. Can you translate that into an equation?

OpenStudy (anonymous):

hmmmm x(x+1)/5? if not that i dont know lol

terenzreignz (terenzreignz):

Okay... relax... ... ... ... again... What is the sum of x and x + 1 ?

OpenStudy (anonymous):

x(x+1)=5x

terenzreignz (terenzreignz):

Not quite... I'm just asking you what the sum of x and x+1 is :D

OpenStudy (anonymous):

idk though lol 5

terenzreignz (terenzreignz):

No... you're complicating things... and that's actually my job :P the sum of x and x+1 is just... <drum roll>

terenzreignz (terenzreignz):

x + x + 1 LOL

OpenStudy (anonymous):

oh my gosh lol im dumb.

terenzreignz (terenzreignz):

Or 2x + 1 Now, what's the difference of x+1 and x?

OpenStudy (anonymous):

the difference between them is 5

terenzreignz (terenzreignz):

Well, the difference between them is (x+1) - x :P

terenzreignz (terenzreignz):

Which is just 1.

OpenStudy (anonymous):

oh ok

terenzreignz (terenzreignz):

Okay, so the sum of these two numbers 2x + 1 is equal to FIVE TIMES THEIR DIFFERENCE. their difference is 1. So... 2x + 1 = 5(1) Now solve :P

OpenStudy (anonymous):

5

terenzreignz (terenzreignz):

lol wut

OpenStudy (anonymous):

so basically 2x+1=5

terenzreignz (terenzreignz):

We're past that now... solve this: 2x + 1 = 5

OpenStudy (anonymous):

x=2

terenzreignz (terenzreignz):

oh yeah.. that

OpenStudy (anonymous):

lol i was working up to it

terenzreignz (terenzreignz):

Okay, great :) So if x = 2 what's the other number (which is exceeds your first number by 1) ?

terenzreignz (terenzreignz):

which exceeds *

OpenStudy (anonymous):

3

terenzreignz (terenzreignz):

So your two numbers are...?

OpenStudy (anonymous):

2,3

terenzreignz (terenzreignz):

Awesome :)

OpenStudy (anonymous):

okay moving on lol

terenzreignz (terenzreignz):

Hey, you're yellow XD

OpenStudy (anonymous):

yay lol